The sum of Father and son’s age is 60 years. Six years ago the age of father is twice the son s age, then the current age of the father is
Answer by richwmiller(17219)

let s= son's age and f= father's age
f+s=60
f=60-s
f-6=2(s-6)
substitute 60-s for f
60-s-6=2(s-6)
combine
54-s=2s-12
add 12
66-s=2s
add s
66=3s
divide by 3
22=s son is 22
f+22=60
subtract 22
f=38 father is 38
check
f-6=2(s-6)
38-6=2(22-6)
32=2*16
32=32
check